### v3

@@ 2 Introduction
Contemporary technologies for recording neural activity can produce multiple time series in each of two or more brain regions simultaneously (e.g., Jun et al., 2017; Steinmetz et al., 2018), enabling identification of cross-regional interactions relevant to cognitive processes that support behavior. One such process is working memory (Miller et al., 2018; Schmidt et al., 2019). During a variety of experimental working memory tasks, strong neural oscillations in the beta range (16-30 Hz) have been observed, leading to the proposal that beta oscillations serve to coordinate activity across regions (Miller et al., 2018). To identify coordinated oscillatory activity, many statistical methods based on coherence, or phase coupling, have been developed, studied, and applied (Belluscio et al., 2012; Klein et al., 2020; Ombao and Pinto, 2022; Urban et al., 2023). As shown in Figure 1, however, oscillatory amplitudes (amplitude envelopes) also vary substantially and may well exhibit statistical dependence across regions. Our aim is to detect whether these slowly-varying amplitude envelopes tend to fluctuate together, and to pinpoint when during a task such coupling occurs. This form of amplitude-based coordination reflects a different aspect of neural communication than phase-based measures and requires statistical tools that can accommodate smooth, non-oscillatory signal components. In this paper, we develop methods for analyzing groups of slowly-varying multiple time series, and apply them to uncover coordinated amplitude activity between brain regions.
@@
A common strategy for quantifying inter-areal association is to compute amplitude–envelope correlations after extracting instantaneous amplitudes via the Hilbert transform, based on region-level summaries of multichannel brain oscillation recordings (Bruns et al., 2000; Zamm et al., 2018; Raghavan et al., 2024). Creating the region-level summaries requires preprossessing, such as by selecting one channel per region (Bruns et al., 2000), applying spatial filters (Zamm et al., 2018), or designating specific channels according to expert judgment (Raghavan et al., 2024). These choices collapse multichannel activity into signals that mix task-relevant and task-irrelevant processes and are vulnerable to spatial leakage and common input, which reduces sensitivity to between-region coupling and obscures the subpopulations that carry the interaction. We instead developed a method to learn low-dimensional and time-localized components directly from multichannel data, without spatial priors. By maximizing between-region dependence (relative to within-region variance) our approach increases sensitivity to inter-areal communication.
@@
The data that motivated this work are local field potentials (LFPs; Buzsáki et al., 2012; Pesaran et al., 2018), recorded from prefrontal cortex (PFC) and visual area V4 during repeated trials of a working memory task (Khanna et al., 2020). PFC is generally considered to exert control aimed at areas involved in perceptual processing (Miller and Cohen, 2001), such as V4, which is active during the retention of higher-order visual information (e.g. color and shape) and attention to visual objects (Orban, 2008; Fries et al., 2001). Both regions exhibit strong oscillatory activity during visual working memory tasks, and their interaction, particularly through oscillatory coupling, has been implicated in supporting working memory function (Sarnthein et al., 1998; Liebe et al., 2012; Miller et al., 2018; Gregoriou et al., 2009).
@@
The data consist of 96 LFP time series for each brain region (192 time series in total), offering millisecond temporal precision, and are relatively precise spatially, as well. These signals reflect neural activity across large populations of neurons within each region and exhibit substantial spatial correlation. To summarize the cross-regional interactions more succinctly, dimensionality reduction becomes a natural and effective step. If we were to analyze the data at a single time point, we would have repeated pairs of 96-dimensional observations, for which canonical correlation analysis (CCA) would be a standard tool to assess statistical dependence. Here we build on a latent variable formulation of CCA known as probabilistic CCA (pCCA), and extend it to handle temporally structured data, as illustrated in Figure 2(b). In the extended model described in Section 3.2, each of the two high-dimensional time series is driven by a latent univariate time series, with the correlation between the two latent time series representing the statistical dependence we seek to identify (see Figure 3). Theorem 4.3 demonstrates that this model yields estimates equivalent to those from a generalization of classical CCA, called multiset CCA (Kettenring, 1971)). This model-based framework allows for flexible modeling of covariance structure and facilitates principled high-dimensional inference.
@@
Related latent variable methods include Gaussian Process Factor Analysis (GPFA, Lakshmanan et al., 2015; Semedo et al., 2020; Yu et al., 2009) and state-space models. GPFA assumes that latent time series follow a low-dimensional Gaussian process, while state-space models typically impose autoregressive dynamics. Although these methods were originally devised to capture shared low-dimensional structure within a single brain region, they can be embedded in a two-step approach to discover cross-regional interaction: (1) apply those methods separately to each brain region, and (2) compute cross-correlations between the resulting latent trajectories. While common in neuroscience, this strategy may lack statistical power when within-region activity does not align with cross-regional communication. For example, in recordings from macaque visual areas V1 and V2, Semedo et al. (2019) found that communication occurred in a dedicated “communication subspace” distinct from the dominant within-region activity. Ignoring this distinction can lead to underestimating cross-regional coupling and reduced sensitivity (Semedo et al., 2020).
@@
To address this, alternative methods have been proposed. Rodu et al. (2018) extended a kernel version of CCA (KCCA; Hardoon et al., 2004) to develop Dynamic Kernel CCA (DKCCA), designed to detect time-varying connectivity between multivariate signals. Gokcen et al. (2022) proposed Delayed Latents Across Groups (DLAG), which extends GPFA to jointly model within- and across-region interactions using structured latent processes. Both methods come with own limitations. For instance, DLAG’s reliance on a structured latent covariance can be limiting for systems with bi-directional coupling via low-dimensional oscillatory dynamics.
@@
In this work, we introduce a more flexible alternative by leaving the latent covariance unspecified, allowing for a fully data-driven estimate of cross-regional correlations. This yields a partial correlation graph representation of neural interactions. Given the high dimensionality of the possible partial correlations across time lags, we adopt a sparse estimation framework, resulting in our method: Latent Dynamic analysis via Sparse banded graphs (LaDynS). While we applied LaDynS to LFP recordings, it could also be used to analyze other slowly-varying multidimensional neural measurements made during repeated-trial cognitive tasks.
@@
In Section 3, after defining the model and proving that it produces a time series generalization of CCA, we give details of the fitting procedure and discuss inference using clustered contiguous significant association based on a de-biased estimate of the precision matrix (Jankova et al., 2015) together with control of false discovery rate (Benjamini and Hochberg, 1995). We also use a state-space formulation to impose a local stationarity assumption (Ombao and Pinto, 2022), pointing out how this can produce a time-varying latent Granger causality assessment. Simulations in Section 4.2 show that our implementation of LaDynS is able to correctly identify the timing of interactions when applied to artificial data designed to be similar to those we analyzed, even in scenarios where existing methods exhibit low statistical power. The simulations support data-analytic results presented in Section 4.4, which show significant cross-regional interaction at roughly $400$ ms after presentation of the stimulus (see Figure 12). This timing coincides with the delay in visual cortical response seen after the onset of a visual stimulus in mice (Chen et al., 2022), humans (Del Cul et al., 2007; Yang et al., 2019), and macaques (Supèr et al., 2001; Schmolesky et al., 1998). The results reinforce the idea that PFC and V4 are involved, together, in working memory. We add discussion in Section 5.
@@ 3.1 Probabilistic CCA for two random vectors
Figure 2: Graphical representation of pCCA models. (a) Model of Bach and Jordan (2005), where $X_{1}$ and $X_{2}$ are random vectors and $Z$ is a random variable. (b) A variation on (a) that facilitates the extension to the case when $X_{1}$ and $X_{2}$ are multivariate time series and $(Z_{1},Z_{2})$ is a bivariate time series.
Given two random vectors $X_{1}\in\mathbb{R}^{d_{1}}$ and $X_{2}\in\mathbb{R}^{d_{2}}$, canonical correlation analysis (CCA) (Hotelling, 1992) finds the sets of weights $w_{1}\in\mathbb{R}^{d_{1}}$ and $w_{2}\in\mathbb{R}^{d_{2}}$ that maximize Pearson’s correlation between linear combinations $w_{1}^{\top}X_{1}$ and $w_{2}^{\top}X_{2}$. This can be rewritten as
@@
where $\mu_{k}\in\mathbb{R}^{d_{k}}$ and $\beta_{k}\in\mathbb{R}^{d_{k}}$ are mean vectors and factor loadings, respectively, and $\epsilon_{k}\overset{\textrm{\tiny{indep}}}{\sim}\operatorname{MVN}(0,\Phi_{k})$ (Bach and Jordan, 2005). Figure 2(a) depicts the dependence of $X_{1}$ and $X_{2}$ on $Z$. The parameters in Eqs. 1 and 2 both yield the same estimate of $\sigma_{cc}$; see Theorem 4.1 for more details.
@@ 3.2 Probabilistic CCA for two time series of random vectors
to highlight the auto-correlations ${\Sigma}_{11}$ and ${\Sigma}_{22}$ within and cross-correlations $\Sigma_{12}$ between the time series, and denote by ${\Sigma}_{kl}^{(t,s)}$, $(t,s)\in[T]^{2}$, the elements of ${\Sigma}_{kl}$. Then ${\Sigma}_{12}^{(t,t)}$ for some fixed $t$ has the same interpretation as $\sigma_{12}$ in Eq. 2. Further, $\Omega_{12}^{(t,s)}/\sqrt{\Omega_{11}^{(t,t)}\Omega_{22}^{(s,s)}}$ is the partial correlation between the two latent time series at times $t$ and $s$. Thus, when an element of $\Omega_{12}$ is non-null, depicted by the red star in the expanded display in Figure 3(b), its coordinates $(t,s)$ and distance (ts)(t-s) from the diagonal indicate at what time in the trial a connectivity happens between two time series, and at what lead or lag, respectively. In our neuroscience application, they represent the timing of connections and direction of information flow between two brain regions. As is common with latent variable models, the signs of the latent covariance $\Sigma$ and therefore the latent precision $\Omega$ are not identifiable. Therefore in the following sections we focus on the estimation and inference for the magnitude of associations, rather than their sign.
As the equivalence between a non-distributional method (CCA) and its probabilistic representation (pCCA), there exists a similar connection between the multiset generalization of CCA introduced by Kettenring (1971) and the dynamic pCCA model in Eqs. 5 and 6. Multiset CCA applied to $2T$ random vectors $\{X_{1}^{(t)},X_{2}^{(t)}:t=1,\dots,T\}$ finds weights $\{w_{1}^{(t)},w_{2}^{(t)}:t=1,\dots,T\}$ that maximize a notion of correlation among linear combinations $\{w_{1}^{(t)\top}X_{1}^{(t)},w_{2}^{(t)\top}X_{2}^{(t)}:t=1,\dots,T\}$. In particular, the GENVAR extension minimizes the generalized variance of these linear combinations, defined as the determinant of their correlation matrix (Wilks, 1932), which we refer to as the canonical correlation matrix:
@@ 3.3.1 Choosing regularization parameters
In graphical LASSO (gLASSO) problems, where the aim is to recover correct partial correlation graphs, penalties are often chosen to minimize the predictive risk (Shao, 1993; Zou et al., 2007; Tibshirani et al., 2012). Our aim is different: only the partial cross-precision matrix $\Omega_{12}$ is of substantive interest, and because minimizing the predictive risk does not select models consistently (Shao, 1993; Zhu and Cribben, 2018) and may thus fail to retrieve non-zero elements of $\Omega_{12}$, we choose instead a value of $\lambda_{\text{cross}}$ that controls the number of false cross-precision discoveries. We proceed by permuting the observed time series in one brain region to create a synthetic dataset that contains no cross-region correlation, then applying LaDynS to that data for a range of values of $\lambda_{\text{cross}}$ and recording the resulting number of significant partial correlation estimates, which are necessarily spurious. We use the smallest $\lambda_{\text{cross}}$ that yields fewer false discoveries than a chosen threshold. We expect this regularization to make similarly few false discoveries on experimental data.
@@ 3.3.2 Fitting LaDynS using coordinate descent
which we can solve efficiently using a number of existing algorithms; here we use the P-gLASSO algorithm of Mazumder and Hastie (2012). Then assuming that all parameters are fixed but a single weight $w_{k}^{(t)}$, Eq. 10 can be re-arranged as the linear problem:
@@
for which an analytical solution is available. That is, our algorithm alternates between updating $\Omega$ and the weights $w_{k}^{(t)}$ until the objective function in Eq. 10 converges. Its computational cost is inexpensive: a single iteration on our cluster server (with $11$ Intel(R) Xeon(R) CPU $2.90$ GHz processors) took in average less than $0.8$ seconds, applied to the experimental data in Section 4.4. A single fit on the same data took $47$ iterations for around $33.57$ seconds until the objective function converged at threshold $0.001$. See Algorithm S1 in Section S2.1 for details and Python package ladyns on github.com/HeejongBong/ladyns.

### v4

@@ 2 Introduction
Contemporary technologies for recording neural activity can produce multiple time series in each of two or more brain regions simultaneously (e.g., jun2017fully; steinmetz2018challenges), enabling identification of cross-regional interactions relevant to cognitive processes that support behavior. One such process is working memory (miller2018wm2; schmidt2019beta). During a variety of experimental working memory tasks, strong neural oscillations in the beta range (16-30 Hz) have been observed, leading to the proposal that beta oscillations serve to coordinate activity across regions (miller2018wm2). To identify coordinated oscillatory activity, many statistical methods based on coherence, or phase coupling, have been developed, studied, and applied (belluscio2012cross; klein2020torus; ombao2022spectral; urban2023oscillating). As shown in Figure 1, however, oscillatory amplitudes (amplitude envelopes) also vary substantially and may well exhibit statistical dependence across regions. Our aim is to detect whether these slowly-varying amplitude envelopes tend to fluctuate together, and to pinpoint when during a task such coupling occurs. This form of amplitude-based coordination reflects a different aspect of neural communication than phase-based measures and requires statistical tools that can accommodate smooth, non-oscillatory signal components. In this paper, we develop methods for analyzing groups of slowly-varying multiple time series, and apply them to uncover coordinated amplitude activity between brain regions.
A common strategy for quantifying inter-areal association is to reduce multichannel recordings within each brain region to a small number of region-level summaries and then compute pairwise association measures between regions. For oscillatory neural signals, this is often done by computing amplitude envelope correlations after extracting instantaneous amplitudes via the Hilbert transform, based on region-level summaries (bruns2000amplitude; zamm2018amplitude; raghavan2024gamma). Constructing these summaries typically requires additional preprocessing, such as selecting a single representative channel per region (bruns2000amplitude), applying spatial filters (zamm2018amplitude), or designating channels based on expert judgment (raghavan2024gamma).
A similar dimension reduction principle underlies methods developed for non-oscillatory recordings such as fMRI. Time Delay (TD) analysis (mitra2015lag; mitra2018spontaneous; raut2019organization) averages voxel-level signals within each region to obtain a single regional time-series. It then compares pairs of regional time=series by measuring how much one signal is delayed relative to the other using cross-covariance, resulting in a single lag value for each region pair that can be further analyzed using PCA. Cyclicity analysis (shahsavarani2020comparing; abraham2024hemodynamic) instead aggregates directionality measures across voxel-level interactions to produce a singed scalar measure that represents region-level summary that captures region-level lead-lag relationships.
@@
These approaches collapse multichannel activity into signals that mix task-relevant and task-irrelevant processes, which reduces sensitivity to between-region coupling and obscures the subpopulations that carry the interaction. We instead developed a method to learn low-dimensional and time-localized components directly from multichannel data, without spatial priors. By maximizing between-region dependence (relative to within-region variance) our approach increases sensitivity to inter-areal communication.
@@
The data that motivated this work are local field potentials (LFPs; buzsaki2012origin; pesaran2018investigating), recorded from prefrontal cortex (PFC) and visual area V4 during repeated trials of a working memory task (khanna2019dynamic). PFC is generally considered to exert control aimed at areas involved in perceptual processing (miller2001integrative), such as V4, which is active during the retention of higher-order visual information (e.g. color and shape) and attention to visual objects (orban2008higher; fries2001modulation). Both regions exhibit strong oscillatory activity during visual working memory tasks, and their interaction, particularly through oscillatory coupling, has been implicated in supporting working memory function (sarnthein1998synchronization; liebe2012theta; miller2018wm2; gregoriou2009high).
@@
The data consist of 96 LFP time series for each brain region (192 time series in total), offering millisecond temporal precision, and are relatively precise spatially, as well. These signals reflect neural activity across large populations of neurons within each region and exhibit substantial spatial correlation. To summarize the cross-regional interactions more succinctly, dimensionality reduction becomes a natural and effective step (gallego2022local). If we were to analyze the data at a single time point, we would have repeated pairs of 96-dimensional observations, for which canonical correlation analysis (CCA) would be a standard tool to assess statistical dependence. Here we build on a latent variable formulation of CCA known as probabilistic CCA (pCCA), and extend it to handle temporally structured data, as illustrated in Figure 2(b). In the extended model described in Section 3.2, each of the two high-dimensional time series is driven by a latent univariate time series, with the correlation between the two latent time series representing the statistical dependence we seek to identify (see Figure 3). Theorem 4.3 demonstrates that this model yields estimates equivalent to those from a generalization of classical CCA, called multiset CCA (kettenring1971canonical)). This model-based framework allows for flexible modeling of covariance structure and facilitates principled high-dimensional inference.
@@
Related latent variable methods include Gaussian Process Factor Analysis (GPFA, lakshmanan2015extracting; semedo2020statistical; yu2009gaussian) and state-space models (goris2014partitioning). GPFA assumes that latent time series follow a low-dimensional Gaussian process, while state-space models typically impose autoregressive dynamics. Although these methods were originally devised to capture shared low-dimensional structure within a single brain region, they can be embedded in a two-step approach to discover cross-regional interaction: (1) apply those methods separately to each brain region, and (2) compute cross-correlations between the resulting latent trajectories. While common in neuroscience, this strategy may lack statistical power when within-region activity does not align with cross-regional communication. For example, in recordings from macaque visual areas V1 and V2, semedo2019cortical found that communication occurred in a dedicated “communication subspace” distinct from the dominant within-region activity. Ignoring this distinction can lead to underestimating cross-regional coupling and reduced sensitivity (semedo2020statistical).
@@
To address this, alternative methods have been proposed. rodu2018detecting extended a kernel version of CCA (KCCA; hardoon2004canonical) to develop Dynamic Kernel CCA (DKCCA), designed to detect time-varying connectivity between multivariate signals. gokcen2021disentangling proposed Delayed Latents Across Groups (DLAG), which extends GPFA to jointly model within- and across-region interactions using structured latent processes. Both methods come with own limitations. For instance, DLAG’s reliance on a structured latent covariance can be limiting for systems with bi-directional coupling via low-dimensional oscillatory dynamics.
@@
In this work, we introduce a more flexible alternative by leaving the latent covariance unspecified, allowing for a fully data-driven estimate of cross-regional correlations. This yields a partial correlation graph representation of neural interactions. Given the high dimensionality of the possible partial correlations across time lags, we adopt a sparse estimation framework (friedman2008sparse), resulting in our method: Latent Dynamic analysis via Sparse banded graphs (LaDynS). While we applied LaDynS to LFP recordings, it could also be used to analyze other slowly-varying multidimensional neural measurements made during repeated-trial cognitive tasks.
@@
In Section 3, after defining the model and proving that it produces a time series generalization of CCA, we give details of the fitting procedure and discuss inference using clustered contiguous significant association based on a de-biased estimate of the precision matrix (jankova2015confidence) together with control of false discovery rate (benjamini1995controlling). We also use a state-space formulation to impose a local stationarity assumption (ombao2022spectral), pointing out how this can produce a time-varying latent Granger causality assessment. Simulations in Section 4.2 show that our implementation of LaDynS is able to correctly identify the timing of interactions when applied to artificial data designed to be similar to those we analyzed, even in scenarios where existing methods exhibit low statistical power. The simulations support data-analytic results presented in Section 4.4, which show significant cross-regional interaction at roughly $400$ ms after presentation of the stimulus (see Figure 12). This timing coincides with the delay in visual cortical response seen after the onset of a visual stimulus in mice (chen2022population), humans (del2007brain; yang2019exploring), and macaques (super2001two; schmolesky1998signal). The results reinforce the idea that PFC and V4 are involved, together, in working memory. We add discussion in Section 5.
@@ 3.1 Probabilistic CCA for two random vectors
Figure 2: Graphical representation of pCCA models. (a) Model of bach2005probabilistic, where $X_{1}$ and $X_{2}$ are random vectors and $Z$ is a random variable. (b) A variation on (a) that facilitates the extension to the case when $X_{1}$ and $X_{2}$ are multivariate time series and $(Z_{1},Z_{2})$ is a bivariate time series.
Given two random vectors $X_{1}\in\mathbb{R}^{d_{1}}$ and $X_{2}\in\mathbb{R}^{d_{2}}$, canonical correlation analysis (CCA) (hotelling1992relations) finds the sets of weights $w_{1}\in\mathbb{R}^{d_{1}}$ and $w_{2}\in\mathbb{R}^{d_{2}}$ that maximize Pearson’s correlation between linear combinations $w_{1}^{\top}X_{1}$ and $w_{2}^{\top}X_{2}$. This can be rewritten as
@@
where $\mu_{k}\in\mathbb{R}^{d_{k}}$ and $\beta_{k}\in\mathbb{R}^{d_{k}}$ are mean vectors and factor loadings, respectively, and $\epsilon_{k}\overset{\textrm{\tiny{indep}}}{\sim}\operatorname{MVN}(0,\Phi_{k})$ (bach2005probabilistic). Figure 2(a) depicts the dependence of $X_{1}$ and $X_{2}$ on $Z$. The parameters in Eqs. 1 and 2 both yield the same estimate of $\sigma_{cc}$; see Theorem 4.1 for more details.
@@ 3.2 Probabilistic CCA for two time series of random vectors
to highlight the auto-correlations ${\Sigma}_{11}$ and ${\Sigma}_{22}$ within and cross-correlations $\Sigma_{12}$ between the time series, and denote by ${\Sigma}_{kl}^{(t,s)}$, $(t,s)\in[T]^{2}$, the elements of ${\Sigma}_{kl}$. Then ${\Sigma}_{12}^{(t,t)}$ for some fixed $t$ has the same interpretation as $\sigma_{12}$ in Eq. 2. Further, $\Omega_{12}^{(t,s)}/\sqrt{\Omega_{11}^{(t,t)}\Omega_{22}^{(s,s)}}$ is the partial correlation between the two latent time series at times $t$ and $s$. Thus, when an element of $\Omega_{12}$ is non-null, depicted by the red star in the expanded display in Figure 3(b), its coordinates $(t,s)$ and distance (ts)(t-s) from the diagonal indicate at what time in the trial a connectivity happens between two time series, and at what lead or lag, respectively. In our neuroscience application, they represent the timing of connections and direction of information flow between two brain regions. As is common with latent variable models, the signs of the latent covariance $\Sigma$ and therefore the latent precision $\Omega$ are not identifiable. This non-identifiability is benign for inference on the directionality in lead-lag connectivity, which is encoded in the support of $\Omega_{12}$, rather than by their signs. (See Figure 3(b).) The directionality remains interpretable even though the signs of the partial-correlations are unidentifiable. Therefore in the following sections we focus on the estimation and inference for the magnitude of associations, rather than their sign.
As the equivalence between a non-distributional method (CCA) and its probabilistic representation (pCCA), there exists a similar connection between the multiset generalization of CCA introduced by kettenring1971canonical and the dynamic pCCA model in Eqs. 5 and 6. Multiset CCA applied to $2T$ random vectors $\{X_{1}^{(t)},X_{2}^{(t)}:t=1,\dots,T\}$ finds weights $\{w_{1}^{(t)},w_{2}^{(t)}:t=1,\dots,T\}$ that maximize a notion of correlation among linear combinations $\{w_{1}^{(t)\top}X_{1}^{(t)},w_{2}^{(t)\top}X_{2}^{(t)}:t=1,\dots,T\}$. In particular, the GENVAR extension minimizes the generalized variance of these linear combinations, defined as the determinant of their correlation matrix (wilks1932certain), which we refer to as the canonical correlation matrix:
@@ 3.3.1 Choosing regularization parameters
In graphical LASSO (gLASSO) problems, where the aim is to recover correct partial correlation graphs, penalties are often chosen to minimize the predictive risk (shao1993linear; zou2007degrees; tibshirani2012degrees). Our aim is different: only the partial cross-precision matrix $\Omega_{12}$ is of substantive interest, and because minimizing the predictive risk does not select models consistently (shao1993linear; zhu2018sparse) and may thus fail to retrieve non-zero elements of $\Omega_{12}$, we choose instead a value of $\lambda_{\text{cross}}$ that controls the number of false cross-precision discoveries. We proceed by permuting the observed time series in one brain region to create a synthetic dataset that contains no cross-region correlation, then applying LaDynS to that data for a range of values of $\lambda_{\text{cross}}$ and recording the resulting number of significant partial correlation estimates, which are necessarily spurious. We use the smallest $\lambda_{\text{cross}}$ that yields fewer false discoveries than a chosen threshold. We expect this regularization to make similarly few false discoveries on experimental data.
@@ 3.3.2 Fitting LaDynS using coordinate descent
which we can solve efficiently using a number of existing algorithms; here we use the P-gLASSO algorithm of mazumder2012graphical. Then assuming that all parameters are fixed but a single weight $w_{k}^{(t)}$, Eq. 10 can be re-arranged as the linear problem:
@@
for which an analytical solution is available. That is, our algorithm alternates between updating $\Omega$ and the weights $w_{k}^{(t)}$ until the objective function in Eq. 10 converges. Its computational cost is inexpensive: a single iteration on our cluster server (with $11$ Intel(R) Xeon(R) CPU $2.90$ GHz processors) took in average less than $0.8$ seconds, applied to the experimental data in Section 4.4. A single fit on the same data took $47$ iterations for around $33.57$ seconds until the objective function converged at threshold $0.001$. Algorithm 1 presents a high level pseudocode sketch of the coordinate descent algorithm. See Section S2.1 for details and Python package ladyns on github.com/HeejongBong/ladyns.
Algorithm 1 Coordinate descent algorithm to fit LaDynS
[…]
0:
$w_{k}^{(t)}\leftarrow\mathbf{1}/\sqrt{\mathbf{1}^{\top}\overline{\operatorname{Var}}[X_{k}^{(t)}]\mathbf{1}}~~\text{and}~~\overline{\Sigma}\leftarrow\overline{\operatorname{Var}}[w_{1}^{(1)\top}X_{1}^{(1)},\dots,w_{2}^{(T)\top}X_{2}^{(T)}].$
1:
[…]
3:  $\Omega\leftarrow\operatorname*{\arg\!\min}_{\Omega^{\prime}}-\log\det(\Omega^{\prime})+\operatorname{tr}(\Omega^{\prime}\overline{\Sigma})+\|\Lambda\odot\Omega^{\prime}\|_{1}$.
4:  for $k$ in 1:2 and $t$ in 1:$T$ do
5:   $w_{k}^{(t)}\leftarrow\operatorname*{\arg\!\min}_{w}{\sum}_{(l,s)\neq(k,t)}w^{\top}\overline{\operatorname{Cov}}\left[X_{k}^{(t)},X_{l}^{(s)}\right]w_{l}^{(s)}\Omega_{kl}^{(t,s)}~~\text{s.t.}~~w^{\top}\overline{\operatorname{Var}}[X_{k}^{(t)}]w=1,$
6:  end for
7:  $\overline{\Sigma}\leftarrow\overline{\operatorname{Var}}\left[w_{1}^{(1)\top}X_{1}^{(1)},\dots,w_{2}^{(T)\top}X_{2}^{(T)}\right].$
8:  if $\max(|\Sigma-\Sigma_{\text{last}}|)<\text{ths}$ then
9:   break
10:  end if
11:end for